ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 90D60, 60C05, 60J20 Campbell, Paul J., Beloit College, U. S. A., and Universität Augsburg, Germany Klondike solitaire Klondike, a playing-card solitaire, may be the world’s most frequently played solitaire game, because of its distribution as software with new computers. We use a state-space Markov-chain approach to analyze simplified models of the game, and we quantify the effect of various rules on the player’s expected score.